OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y =1/2 x + 3. y = -2x - 7 y = -2x +5 y = 2x + 5 y = 2x - 7

OpenStudy (anonymous):

y =1/2 x + 3. m = 1/2 perpendicular = - 1/m = -2 y =-2 x + b. now you can choose y = -2x - 7 y = -2x +5 y =-2 x + b. (-3, -1) -1=-2*(-3)+b -1 = 6 + b b = -7 y = -2x - 7

OpenStudy (anonymous):

don't just give the answer, walk them through the problem